A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. ​n=11 p=0.1 x<4
The probability of x<4 successes is
​(Round to four decimal places as​ needed.)

Answer by Boreal(15235)

This is p x=0,1,2,3
x=0 has prob. .9^11=0.3138
x=1 has prob. 11*0.9^10*0.1=0.3835
x=2 has prob. 11C2*0.9^9*0.1^2=0.2131
x=3 has prob. 11C3*09.^8*0.1^3=0.0710
that prob. sum is 0.9815 when rounding is considered
calculator 2nd VARS binomcdf(11,0.1,3) not 4 because it is < 4, not <=4.